Solve by elimination method.
Jill is 27.5 kilometers away from Joe. Both begin to walk toward each other at the same time. Jill walks at 3 kilometers per hour. They meet in 5 hours. How fast is Joe walking?

Found 3 solutions by ankor@dixie-net.com, MathTherapy, ikleyn:
Answer by ankor@dixie-net.com(22740)

Jill is 27.5 kilometers away from Joe.
Both begin to walk toward each other at the same time.
Jill walks at 3 kilometers per hour.
They meet in 5 hours.
How fast is Joe walking?
:
let j = Joe's walking speed
:
Why would you use elimination, there is only one unknown?
Toward each other, their speed are added
write a dist equation, dist = time * speed
5(j + 3) = 27.5
5j + 15 = 27.5
5j = 27.5 - 15
5j = 12.5
j = 12.5/5
j = 2.5 km/hr is Joe's speed

Answer by MathTherapy(10552)


Solve by elimination method.
Jill is 27.5 kilometers away from Joe. Both begin to walk toward each other at the same time. Jill walks at 3 kilometers per hour. They meet in 5 hours. How fast is Joe walking?
Let Joe's speed be S
Then we get: 5S + 5(3) = 27.5
5S + 15 = 27.5
5S = 12.5
Joe's speed or

Answer by ikleyn(52778)

.

The instruction in the post  "Solve by elimination method"  is  IRRELEVANT  to the problem.


I am 129% sure,  that this instruction came by mistake from another problem.